15-th Conference of the International Linear Algebra Society

March 27, 2008

Professor Dr. Zeyad Abdel Aziz Al Zhour

Department of Mathematics Zarqa Private University Zarqa , Jordan

Dear Professor Al Zhour:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results in the Contributed Papers Session.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008. Some useful information is available at the conference web-page

http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4653 e-mail: verde@star.izt.uam.mx